OpenStudy (anonymous):

Find the slope intercept form for the line satisfying the following conditions. x-intercept 3, y-intercept 1/5

OpenStudy (anonymous):

First Write the Equarion of the line using intercept form

OpenStudy (anonymous):

x/a + y/b = 1 a = x intercept b = y intercept

OpenStudy (anonymous):

(3, 0) x-intercept (0, 1/5) y-intercept find the slope from these points slope = m = (1/5 - 0)/(0-3) m = (1/5)/-3 =-1/15 now slope = -1/15 and y-intercept = 1/5 can you write the equation now???

OpenStudy (anonymous):

Equn of intercept from line is given by x/a +y/b=1 so you have given a=3 y=1/5 so easily calculate x/3+5y=1